OpenStudy (anonymous):

What is true about the solutions of a quadratic equation when the radicand in the quadratic formula is negative? No real solutions Two identical rational solutions Two different rational solutions Two irrational solutions

OpenStudy (lgbasallote):

hint: b^2 -4ac > 0 means 2 different solution b^2 -4ac = 0 means 2 identical solutions b^2 - 4ac < 0 means no real solutions does that help?

OpenStudy (anonymous):

a little, but what is a radicand?

OpenStudy (lgbasallote):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] the radicand is the expression *in* the square root..in other worrds b^2 -4ac

OpenStudy (anonymous):

ohhh okay thanks!

OpenStudy (lgbasallote):

welcome

OpenStudy (anonymous):

so...would it be two different rational

OpenStudy (lgbasallote):

why do you say that?

OpenStudy (lgbasallote):

if the radicand is negative...then is it greater than 0?equal to 0? or less than 0?

OpenStudy (anonymous):

less than zero

OpenStudy (anonymous):

ohhh aha forgot about your hint

OpenStudy (lgbasallote):

haha lol =))

OpenStudy (anonymous):

Whats the answer???????????